U={1,2,3,4,5,6,7,8,9,10}



A={2,3,4,5,8}



B={2,4,6,8,10}



C={1,3,5,7,9}




Instruction: Solve the given below using the sets above



1.( (C U A)-B)’ ∩ (B’ U A)



2.(((BUA)’ U (CUA)’) U ((A-B)’ ∩ (B-C)’))

1
Expert's answer
2022-05-09T11:12:43-0400

"2.(BUA)=\{2,3,4,5,6,8,10\}\\(BUA)'=\{1,7,9\}"

"(CUA)=\{1,2,3,4,5,7,8,9\}\\(CUA)'=\{6,10\}"

"((BUA)’ U (CUA)’)=\{1,6,7,9,10\}"

"A-B=\{3,5\}"

"(A-B)'=\{1,2,4,6,7,8,9,10\}"

"B-C=\{2,4,6,8,10\}"

"(B-C)'=\{1,3,5,7,9\}"

"(A-B)' \cap(B-C)'=\{1,7,9\}"

"(((BUA)’ U (CUA)’) U ((A-B)’ ∩ (B-C)’))=\{1,6,7,9,10\}"

1. "(CUA)=\{1,2,3,4,5,7,8,9\}"

"(CUA)-B=\{1,3,5,7,9\}"

"((CUA)-B)'=\{2,4,6,8,10\}"

"B'=\{1,3,5,7,9\}"

"B' U A=\{1,2,3,4,5,7,8,9\}"

"( (C U A)-B)’ ∩ (B’ U A)=\{2,4,8\}"